# Stochastic Optimal Control of an Industrial Power-to-Heat System with High-Temperature Heat Pump and Thermal Energy Storage

### v2

@@ F.2 Estimation of Ornstein-Uhlenbeck Parameters
In roussas2003introduction a derivation for the maximum-likelihood estimator of a bivariate Gaussian process can be found. Although our model differs and extends this approach, we use a similar approach to determine the estimator. The maximum-likelihood estimator Holy2018 of the parameter $\theta$ are obtained by maximizing the log-likelihood function
@@
The necessary condition $\nabla_{\theta}\ell(\theta)=0$ yields a potential maximizer of the log-likelihood function. First we are going to calculate $\nabla_{\theta}A_{n}(\theta)$. The partial derivatives are given by
@@
$\frac{\partial\,A_{n}(\theta)}{\partial p_{W}}$ $=\frac{2}{(1-\rho^{2})}\bigg{[}-\frac{y_{n-1}^{W}}{\Sigma_{W}}\frac{(y^{W}-p_{W}y_{n-1}^{W})}{\Sigma_{W}}+\rho\frac{y_{n-1}^{W}}{\Sigma_{W}}\frac{(y^{S}-p_{S}y_{n-1}^{S}-q_{S}y_{n-1}^{W})}{\Sigma_{S}}\bigg{]}$ (159)
$\frac{\partial\,A_{n}(\theta)}{\partial p_{S}}$ $=\frac{2}{(1-\rho^{2})}\bigg{[}-\frac{y_{n-1}^{S}}{\Sigma_{S}}\frac{(y^{S}-p_{S}y_{n-1}^{S}-q_{S}y_{n-1}^{W})}{\Sigma_{S}}+\rho\frac{y_{n-1}^{S}}{\Sigma_{S}}\frac{(y^{W}-p_{W}y_{n-1}^{W})}{\Sigma_{W}}\bigg{]}$ (160)
$\frac{\partial\,A_{n}(\theta)}{\partial q_{S}}$ $=\frac{2}{(1-\rho^{2})}\bigg{[}-\frac{y_{n-1}^{W}}{\Sigma_{S}}\frac{(y^{S}-p_{S}y_{n-1}^{S}-q_{S}y_{n-1}^{W})}{\Sigma_{S}}+\rho\frac{y_{n-1}^{W}}{\Sigma_{S}}\frac{(y^{W}-p_{W}y_{n-1}^{W})}{\Sigma_{W}}\bigg{]}$ (161)
$\frac{\partial\,A_{n}(\theta)}{\partial\Sigma_{W}^{2}}$ $=\frac{1}{(1-\rho^{2})}\bigg{[}-\frac{(y^{W}-p_{W}y_{n-1}^{W})^{2}}{\Sigma_{W}^{4}}+\rho\frac{(y^{W}-p_{W}y_{n-1}^{W})(y^{S}-p_{S}y_{n-1}^{S}-q_{S}y_{n-1}^{W})}{\Sigma_{W}^{3}\Sigma_{S}}\bigg{]}$ (162)
$\frac{\partial\,A_{n}(\theta)}{\partial\Sigma_{S}^{2}}$ $=\frac{1}{(1-\rho^{2})}\bigg{[}-\frac{(y^{S}-p_{S}y_{n-1}^{S}-q_{S}y_{n-1}^{W})^{2}}{\Sigma_{S}^{4}}+\rho\frac{(y^{W}-p_{W}y_{n-1}^{W})(y^{S}-p_{S}y_{n-1}^{S}-q_{S}y_{n-1}^{W})}{\Sigma_{W}\Sigma_{S}^{3}}\bigg{]}$ (163)
and lastly
$\frac{\partial\,A_{n}(\theta)}{\partial\rho}=\frac{2}{(1-\rho^{2})^{2}}\bigg{[}\rho\bigg{(}\frac{(y^{W}-p_{W}y_{n-1}^{W})^{2}}{\Sigma_{W}^{2}}$ $+\frac{(y^{S}-p_{S}y_{n-1}^{S}-q_{S}y_{n-1}^{W})^{2}}{\Sigma_{S}^{2}}$ (164)
$-(1+\rho^{2})\frac{(y^{W}-p_{W}y_{n-1}^{W})(y^{S}-p_{S}y_{n-1}^{S}-q_{S}y_{n-1}^{W})}{\Sigma_{W}\Sigma_{S}}\bigg{)}\bigg{]}.$ (165)
Now this leads to the partial derivatives for $\ell(\theta)$
$\frac{\partial\,\ell(\theta)}{\partial p_{W}}$ $=-\frac{1}{2}\sum\limits_{n=1}^{N}\frac{\partial\,A_{n}(\theta)}{\partial p_{W}},\quad\frac{\partial\,\ell(\theta)}{\partial p_{S}}=-\frac{1}{2}\sum\limits_{n=1}^{N}\frac{\partial\,A_{n}(\theta)}{\partial p_{S}},\quad\frac{\partial\,\ell(\theta)}{\partial q_{S}}=-\frac{1}{2}\sum\limits_{n=1}^{N}\frac{\partial\,A_{n}(\theta)}{\partial q_{S}}$ (166)
@@
$\frac{\partial\,\ell(\theta)}{\partial\Sigma_{W}^{2}}$ $=-\frac{N}{2\Sigma_{W}^{2}}-\frac{1}{2}\sum\limits_{n=1}^{N}\frac{\partial\,A_{n}(\theta)}{\partial\Sigma_{W}^{2}},\quad\frac{\partial\,\ell(\theta)}{\partial\Sigma_{S}^{2}}=-\frac{N}{2\Sigma_{S}^{2}}-\frac{1}{2}\sum\limits_{n=1}^{N}\frac{\partial\,A_{n}(\theta)}{\partial\Sigma_{S}^{2}}$ (167)
$\frac{\partial\,\ell(\theta)}{\partial\rho}$ $=\frac{N\rho}{(1-\rho^{2})}-\frac{1}{2}\sum\limits_{n=1}^{N}\frac{\partial\,A_{n}(\theta)}{\partial\rho}.$ (168)
@@
We will use the following notation in order to provide a compact derivation of the maximum-likelihood estimators of $\theta$. For that let
$A$ $=\sum\limits_{n=1}^{N}y_{n}^{W}y_{n-1}^{W},~{}B=\sum\limits_{n=1}^{N}(y_{n-1}^{W})^{2},~{}C=\sum\limits_{n=1}^{N}y_{n-1}^{S}y_{n}^{S},~{}D=\sum\limits_{n=1}^{N}(y_{n-1}^{S})^{2},~{}E=\sum\limits_{n=1}^{N}y_{n-1}^{W}y_{n}^{S},$ (169)
$F$ $=\sum\limits_{n=1}^{N}y_{n-1}^{S}y_{n}^{W},~{}G=\sum\limits_{n=1}^{N}y_{n-1}^{W}y_{n-1}^{S},~{}H=\sum\limits_{n=1}^{N}(y_{n}^{W}-p_{W}y_{n-1}^{W})^{2},$ (170)
$I$ $=\sum\limits_{n=1}^{N}(y_{n}^{S}-p_{S}y_{n-1}^{S}-q_{S}y_{n-1}^{W})^{2},~{}J=\sum\limits_{n=1}^{N}(y_{n}^{W}-p_{W}y_{n-1}^{W})(y_{n}^{S}-p_{S}y_{n-1}^{S}-q_{S}y_{n-1}^{W}).$ (171)
@@
$\frac{\partial\,\ell(\theta)}{\partial p_{W}}$ $=\frac{1}{\Sigma_{W}^{2}}(A-p_{W}B)-\frac{\rho}{\Sigma_{W}\Sigma_{S}}(E-p_{S}G-q_{S}B)=0,$ (172)
$\frac{\partial\,\ell(\theta)}{\partial p_{S}}$ $=\frac{1}{\Sigma_{S}^{2}}(C-p_{S}D-q_{S}G)-\frac{\rho}{\Sigma_{W}\Sigma_{S}}(F-p_{W}G)=0,$ (173)
$\frac{\partial\,\ell(\theta)}{\partial q_{S}}$ $=\frac{1}{\Sigma_{S}^{2}}(E-p_{S}G-q_{S}B)-\frac{\rho}{\Sigma_{W}\Sigma_{S}}(A-p_{W}B)=0.$ (174)
Multiplying the first equation by $\frac{\rho\Sigma_{W}}{\Sigma_{W}\Sigma_{S}}$ and subtracting the third we get
$-\frac{\rho^{2}}{\Sigma_{S}^{2}}(E-p_{S}G-q_{S}B)-\frac{1}{\Sigma_{S}^{2}}(E-p_{S}G-q_{S}B)=0~{}\Longleftrightarrow~{}\bigg{(}\frac{\rho^{2}}{\Sigma_{S}^{2}}-\frac{1}{\Sigma_{S}^{2}}\bigg{)}(E-p_{S}G-q_{S}B)=0.$ (175)
From this we can deduce that the first and or second factor needs to be zero. If the first factor is zero, then $\rho^{2}=1$ which by assumption is excluded. Therefore, $E-p_{S}G-q_{S}B=0$, which yields the estimator
@@
Rearranging the second and third equation for $\frac{\rho}{\Sigma_{W}\Sigma_{S}}$ yields
@@
Using equation (177) leads to an expression for $p_{S}$ given according to
@@
Next setting the derivatives with respect to $\Sigma_{W}^{2},\Sigma_{S}^{2}$ and $\rho$ to zero
$\frac{\partial\,\ell(\theta)}{\partial\Sigma_{W}^{2}}$ $=-\frac{N}{2\Sigma_{W}^{2}}+\frac{H}{2(1-\rho^{2})\Sigma_{W}^{4}}-\frac{\rho J}{2(1-\rho^{2})\Sigma_{W}\Sigma_{S}^{3}}=0,$ (180)
$\frac{\partial\,\ell(\theta)}{\partial\Sigma_{S}^{2}}$ $=-\frac{N}{2\Sigma_{S}^{2}}+\frac{I}{2(1-\rho^{2})\Sigma_{S}^{4}}-\frac{\rho J}{2(1-\rho^{2})\Sigma_{W}^{3}\Sigma_{S}}=0,$ (181)
$\frac{\partial\,\ell(\theta)}{\partial\rho}$ $=\frac{N\rho}{1-\rho^{2}}-\frac{1}{(1-\rho^{2})^{2}}\bigg{(}\rho\bigg{[}\frac{H}{\Sigma_{W}^{2}}+\frac{I}{\Sigma_{S}^{2}}\bigg{]}-\frac{1+\rho^{2}}{\Sigma_{W}\Sigma_{S}}J\bigg{)}=0,$ (182)
Rearranging with respect to $1-\rho^{2}$ yields
@@
From the second equalitiy we conclude that $\widehat{\Sigma}_{W}^{2}=\frac{\Sigma_{S}^{2}H}{I}$. Plugging this into the third equality we get
@@
Using this expression in the first equality in (184) we get
@@
$\widehat{\Sigma}_{S}^{2}=\frac{1}{N}I=\frac{1}{N}\sum\limits_{n=1}^{N}(y_{n}^{S}-\widehat{p}_{S}y_{n-1}^{S}-\widehat{q}_{S}y_{n-1}^{W})^{2}\text{ and }\widehat{\Sigma}_{W}^{2}=\frac{1}{N}H=\frac{1}{N}\sum\limits_{n=1}^{N}(y_{n}^{W}-\widehat{p}_{W}y_{n-1}^{W})^{2}.$ (187)
Moreover, we get that $\widehat{\rho}=\frac{\widehat{\Sigma}_{WS}}{{\widehat{\Sigma}_{W}}{\widehat{\Sigma}_{S}}}$ with the maximum-likelihood estimate of the covariance
@@
By using the estimators for $\widehat{\theta}=(\widehat{p}_{W},\widehat{p}_{S},\widehat{q}_{S},\widehat{\Sigma}_{W}^{2},\widehat{\Sigma}_{S}^{2},\widehat{\rho})$ and (154) we obtain estimators
$\widehat{\lambda}_{W}=-\frac{\log(\widehat{p}_{W})}{\Delta t},\qquad\widehat{\sigma}_{W}^{2}=\frac{2\widehat{\lambda}_{W}\widehat{\Sigma}_{W}^{2}}{1-e^{-2\widehat{\lambda}_{W}\Delta t}},$ (189)
for the parameters of the Ornstein-Uhlenbeck process $Y^{W}$ and for the ones of $Y^{S}$
@@
where $\widehat{\lambda}_{W},\widehat{\sigma}_{W}^{2}$ are the maximum-likelihood estimators for the parameters of $Y^{W}$ and
@@
F.3 Wind Turbine Power
The power output of the WT $P^{\text{W}}$ induced by a specific wind speed $W$ can be calculated with its corresponding power curve $P_{\text{WT}}(\cdot)$ via $P^{\text{W}}=P_{\text{WT}}(W)$. In this work, a Vestas V150-4.2  [^3] Vestas, https://www.vestas.com/en/products/4-mw-platform/V150-4-2-MW is used for numerical experiments. The corresponding power curve data is shown in Figure 10, which is available in a database.  [^4] The wind power, www.thewindpower.net/turbine_en_1490_vestas_v150-4000-4200.php
Figure 10: Power curve of a Vestas V150-4.2 with given data (blue bars) and its different operation zones. In region 1 no energy is produced until the cut-in speed of 3 m/s is reached. After that, in region 2 the energy production increases until a certain threshold speed is met. From there, the energy level stays constant at the rated power generation of 4200 kW in region 3. If the wind speed reaches the cut-out speed of 22.5 m/s, no energy is generated in region 4 in order to prevent damage on the WT.
The data of the power curve can be separated into four different regions and suggest a segmented modeling approach for the wind power curve. The regions are classified as follows: Low wind speed where no energy is produced until the cut-in speed of $w_{in}=3$ m/s is met, denoted as region 1 ($R_{1}=[0,w_{in})$). From cut-in to a rated level of $w_{r}=11.5$ m/s the WT power generation increases with the underlying wind speed in region 2 ($R_{2}=[w_{in},w_{\text{Pmax}})$), until reaching its rated power generation of $P^{\text{W}}_{\max}=4200$ kW. In region 3 ($R_{3}=[w_{\text{Pmax}},w_{out})$) the power output stays constant on the rated maximal outout. In order to prevent damage on the WT for high speeds, the generation will turn off at a cut-out speed of $w_{out}=22.5$ m/s summarized in region 4 ($R_{4}=[w_{out},\infty$). Note that the power curve takes constant values except for region 2, where data has to be approximated by a function $P^{\text{W}}_{R_{2}}$. Moreover, we choose a polynomial of degree 6 WANG2019109422 , to fit the underlying data points of the power curve in region 2 see Figure 11, i.e.
$P^{\text{W}}_{R_{2}}(w)=a_{6}w^{6}+a_{5}w^{5}+a_{3}w^{4}+a_{3}w^{3}+a_{2}w^{2}+a_{1}w+a_{0}$ (192)
with coefficients $a_{0},a_{1},a_{2},a_{3},a_{4},a_{5},a_{6}\in\mathbb{R}$ which can be found in Table 3. We are aware that the behavior in region 2 can be mimicked by a cubic polynomial WANG2019109422 . However, our motivation for a polynomial of degree 6 is to ensure a smooth transition from the power output in region 2 to the rated power in region 3.

### v3

@@ G.2 Estimation of Ornstein-Uhlenbeck Parameters
In roussas2003introduction a derivation for the maximum-likelihood estimator Holy2018 of a bivariate Gaussian process can be found. Although our model differs and extends this approach, we use a similar approach to determine the estimator. The maximum-likelihood estimator of the parameter $\theta$ are obtained by maximizing the log-likelihood function
@@
The necessary condition $\nabla_{\theta}\ell(\theta)=0$, yields a potential maximizer of the log-likelihood function. First, we are going to calculate $\nabla_{\theta}A_{n}(\theta)$. The partial derivatives are given by
$\begin{split}\frac{\partial\,A_{n}(\theta)}{\partial p_{W}}&=q_{2}\bigg[-\frac{y_{n-1}^{W}}{\Sigma_{W}}\frac{(y^{W}-p_{W}y_{n-1}^{W})}{\Sigma_{W}}+\rho\frac{y_{n-1}^{W}}{\Sigma_{W}}\frac{(y^{S}-p_{S}y_{n-1}^{S}-q_{S}y_{n-1}^{W})}{\Sigma_{S}}\bigg],\\ \frac{\partial\,A_{n}(\theta)}{\partial p_{S}}&=q_{2}\bigg[-\frac{y_{n-1}^{S}}{\Sigma_{S}}\frac{(y^{S}-p_{S}y_{n-1}^{S}-q_{S}y_{n-1}^{W})}{\Sigma_{S}}+\rho\frac{y_{n-1}^{S}}{\Sigma_{S}}\frac{(y^{W}-p_{W}y_{n-1}^{W})}{\Sigma_{W}}\bigg],\\ \frac{\partial\,A_{n}(\theta)}{\partial q_{S}}&=q_{2}\bigg[-\frac{y_{n-1}^{W}}{\Sigma_{S}}\frac{(y^{S}-p_{S}y_{n-1}^{S}-q_{S}y_{n-1}^{W})}{\Sigma_{S}}+\rho\frac{y_{n-1}^{W}}{\Sigma_{S}}\frac{(y^{W}-p_{W}y_{n-1}^{W})}{\Sigma_{W}}\bigg],\\ \frac{\partial\,A_{n}(\theta)}{\partial\Sigma_{W}^{2}}&=q_{1}\bigg[-\frac{(y^{W}-p_{W}y_{n-1}^{W})^{2}}{\Sigma_{W}^{4}}+\rho\frac{(y^{W}-p_{W}y_{n-1}^{W})(y^{S}-p_{S}y_{n-1}^{S}-q_{S}y_{n-1}^{W})}{\Sigma_{W}^{3}\Sigma_{S}}\bigg],\\ \frac{\partial\,A_{n}(\theta)}{\partial\Sigma_{S}^{2}}&=q_{1}\bigg[-\frac{(y^{S}-p_{S}y_{n-1}^{S}-q_{S}y_{n-1}^{W})^{2}}{\Sigma_{S}^{4}}+\rho\frac{(y^{W}-p_{W}y_{n-1}^{W})(y^{S}-p_{S}y_{n-1}^{S}-q_{S}y_{n-1}^{W})}{\Sigma_{W}\Sigma_{S}^{3}}\bigg],\end{split}$ (138)
with $q_{1}=\frac{1}{(1-\rho^{2})},~q_{2}=\frac{2}{(1-\rho^{2})}$ and finally
@@
$\begin{split}\frac{\partial\,A_{n}(\theta)}{\partial\rho}&=\frac{2}{(1-\rho^{2})^{2}}\bigg[\rho\bigg(\frac{(y^{W}-p_{W}y_{n-1}^{W})^{2}}{\Sigma_{W}^{2}}+\frac{(y^{S}-p_{S}y_{n-1}^{S}-q_{S}y_{n-1}^{W})^{2}}{\Sigma_{S}^{2}}\\ &\hskip 8.5359pt-(1+\rho^{2})\frac{(y^{W}-p_{W}y_{n-1}^{W})(y^{S}-p_{S}y_{n-1}^{S}-q_{S}y_{n-1}^{W})}{\Sigma_{W}\Sigma_{S}}\bigg)\bigg].\end{split}$ (139)
This leads to the partial derivatives for $\ell(\theta)$
@@
$\begin{split}\frac{\partial\,\ell(\theta)}{\partial p_{W}}&=-\frac{1}{2}\sum\limits_{n=1}^{N}\frac{\partial\,A_{n}(\theta)}{\partial p_{W}},\quad\frac{\partial\,\ell(\theta)}{\partial p_{S}}=-\frac{1}{2}\sum\limits_{n=1}^{N}\frac{\partial\,A_{n}(\theta)}{\partial p_{S}},\quad\frac{\partial\,\ell(\theta)}{\partial q_{S}}=-\frac{1}{2}\sum\limits_{n=1}^{N}\frac{\partial\,A_{n}(\theta)}{\partial q_{S}}\\ \frac{\partial\,\ell(\theta)}{\partial\Sigma_{W}^{2}}&=-\frac{N}{2\Sigma_{W}^{2}}-\frac{1}{2}\sum\limits_{n=1}^{N}\frac{\partial\,A_{n}(\theta)}{\partial\Sigma_{W}^{2}},\quad\frac{\partial\,\ell(\theta)}{\partial\Sigma_{S}^{2}}=-\frac{N}{2\Sigma_{S}^{2}}-\frac{1}{2}\sum\limits_{n=1}^{N}\frac{\partial\,A_{n}(\theta)}{\partial\Sigma_{S}^{2}}\\ \frac{\partial\,\ell(\theta)}{\partial\rho}&=\frac{N\rho}{(1-\rho^{2})}-\frac{1}{2}\sum\limits_{n=1}^{N}\frac{\partial\,A_{n}(\theta)}{\partial\rho}.\end{split}$ (140)
@@
We use the following notation to provide a compact derivation of the maximum-likelihood estimator of $\theta$, for this let
$\begin{split}A&=\sum\limits_{n=1}^{N}y_{n}^{W}y_{n-1}^{W},~B=\sum\limits_{n=1}^{N}(y_{n-1}^{W})^{2},~C=\sum\limits_{n=1}^{N}y_{n-1}^{S}y_{n}^{S},~D=\sum\limits_{n=1}^{N}(y_{n-1}^{S})^{2},~E=\sum\limits_{n=1}^{N}y_{n-1}^{W}y_{n}^{S},\\ F&=\sum\limits_{n=1}^{N}y_{n-1}^{S}y_{n}^{W},~G=\sum\limits_{n=1}^{N}y_{n-1}^{W}y_{n-1}^{S},~H=\sum\limits_{n=1}^{N}(y_{n}^{W}-p_{W}y_{n-1}^{W})^{2},\\ I&=\sum\limits_{n=1}^{N}(y_{n}^{S}-p_{S}y_{n-1}^{S}-q_{S}y_{n-1}^{W})^{2},~J=\sum\limits_{n=1}^{N}(y_{n}^{W}-p_{W}y_{n-1}^{W})(y_{n}^{S}-p_{S}y_{n-1}^{S}-q_{S}y_{n-1}^{W}).\end{split}$ (141)
@@
$\begin{split}\frac{\partial\,\ell(\theta)}{\partial p_{W}}&=\frac{1}{\Sigma_{W}^{2}}(A-p_{W}B)-\frac{\rho}{\Sigma_{W}\Sigma_{S}}(E-p_{S}G-q_{S}B)=0,\\ \frac{\partial\,\ell(\theta)}{\partial p_{S}}&=\frac{1}{\Sigma_{S}^{2}}(C-p_{S}D-q_{S}G)-\frac{\rho}{\Sigma_{W}\Sigma_{S}}(F-p_{W}G)=0,\\ \frac{\partial\,\ell(\theta)}{\partial q_{S}}&=\frac{1}{\Sigma_{S}^{2}}(E-p_{S}G-q_{S}B)-\frac{\rho}{\Sigma_{W}\Sigma_{S}}(A-p_{W}B)=0.\end{split}$ (142)
Multiplying the first equation by $\frac{\rho\Sigma_{W}}{\Sigma_{W}\Sigma_{S}}$ and subtracting the third, yields
$\begin{split}&-\frac{\rho^{2}}{\Sigma_{S}^{2}}(E-p_{S}G-q_{S}B)-\frac{1}{\Sigma_{S}^{2}}(E-p_{S}G-q_{S}B)=0\\ \Longleftrightarrow~&\bigg(\frac{\rho^{2}}{\Sigma_{S}^{2}}-\frac{1}{\Sigma_{S}^{2}}\bigg)(E-p_{S}G-q_{S}B)=0.\end{split}$ (143)
From this, we can deduce that the first and/or second factor needs to be zero. If the first factor is zero, then $\rho^{2}=1$, which by assumption is excluded. Therefore, $E-p_{S}G-q_{S}B=0$ holds, which gives the estimator
@@
Rearranging the second and third equation for $\frac{\rho}{\Sigma_{W}\Sigma_{S}}$ takes the form
@@
Using (145), leads to an expression for $p_{S}$ given according to
@@
Next, setting the derivatives with respect to $\Sigma_{W}^{2},\Sigma_{S}^{2}$ and $\rho$ to zero
$\begin{split}\frac{\partial\,\ell(\theta)}{\partial\Sigma_{W}^{2}}&=-\frac{N}{2\Sigma_{W}^{2}}+\frac{H}{2(1-\rho^{2})\Sigma_{W}^{4}}-\frac{\rho J}{2(1-\rho^{2})\Sigma_{W}\Sigma_{S}^{3}}=0,\\ \frac{\partial\,\ell(\theta)}{\partial\Sigma_{S}^{2}}&=-\frac{N}{2\Sigma_{S}^{2}}+\frac{I}{2(1-\rho^{2})\Sigma_{S}^{4}}-\frac{\rho J}{2(1-\rho^{2})\Sigma_{W}^{3}\Sigma_{S}}=0,\\ \frac{\partial\,\ell(\theta)}{\partial\rho}&=\frac{N\rho}{1-\rho^{2}}-\frac{1}{(1-\rho^{2})^{2}}\bigg(\rho\bigg[\frac{H}{\Sigma_{W}^{2}}+\frac{I}{\Sigma_{S}^{2}}\bigg]-\frac{1+\rho^{2}}{\Sigma_{W}\Sigma_{S}}J\bigg)=0,\end{split}$ (148)
and rearranging with respect to $1-\rho^{2}$ leading to
@@
From the second equality, we conclude that $\widehat{\Sigma}_{W}^{2}=\frac{\Sigma_{S}^{2}H}{I}$. Plugging this into the third equality, gives
@@
Using $\widehat{\rho}$ in the first equality in (150) implies
@@
$\widehat{\Sigma}_{S}^{2}=\frac{I}{N}=\frac{1}{N}\sum\limits_{n=1}^{N}(y_{n}^{S}-\widehat{p}_{S}y_{n-1}^{S}-\widehat{q}_{S}y_{n-1}^{W})^{2}\quad\text{and}\quad\widehat{\Sigma}_{W}^{2}=\frac{H}{N}=\frac{1}{N}\sum\limits_{n=1}^{N}(y_{n}^{W}-\widehat{p}_{W}y_{n-1}^{W})^{2}.$ (153)
Moreover, we get $\widehat{\rho}=\frac{\widehat{\Sigma}_{WS}}{{\widehat{\Sigma}_{W}}{\widehat{\Sigma}_{S}}}$, with the maximum-likelihood estimate of the covariance
@@
[…]
$\widehat{\lambda}_{W}=-\frac{\log(\widehat{p}_{W})}{\Delta t},\quad\widehat{\sigma}_{W}^{2}=\frac{2\widehat{\lambda}_{W}\widehat{\Sigma}_{W}^{2}}{1-e^{-2\widehat{\lambda}_{W}\Delta t}},$ (155)
for the parameters of the process $Y^{W}$ and for the process $Y^{S}$
@@
where $\widehat{\lambda}_{W},\widehat{\sigma}_{W}^{2}$ are the maximum-likelihood of $Y^{W}$ and where
@@
G.3 Wind Turbine Power
The WT power output $P^{\text{W}}$ induced by a specific wind speed $W$ can be calculated with its corresponding power curve $P_{\text{WT}}(\cdot)$, via $P^{\text{W}}=P_{\text{WT}}(W)$. In this work, a Vestas V150-4.2333Vestas, https://www.vestas.com/en/products/4-mw-platform/V150-4-2-MW is used for numerical experiments. The corresponding power curve data is shown in Figure 10, which is available in a database [^4] The wind power, www.thewindpower.net/turbine_en_1490_vestas_v150-4000-4200.php.
Figure 10: Power curve of a Vestas V150-4.2 with given data (blue bars) and its different operation zones. In region 1 no energy is produced until the cut-in speed of 3 m/s is reached. After that, in region 2 the energy production increases until a certain threshold speed is met. From there, the energy level stays constant at the rated power generation of 4200 kW,  in region 3. If the wind speed reaches the cut-out speed of 22.5 m/s, no energy is generated in region 4 in order to prevent damage on the WT.
The data of the power curve can be separated into four different regions and suggest a segmented modeling approach for the wind power curve. The regions are classified as follows: Low wind speed, where no energy is produced until the cut-in speed of $w_{\text{in}}=3$ m/s is met, denoted as region 1 ($R_{1}=[0,w_{\text{in}})$). From cut-in to a rated level of $w_{\text{r}}=11.5$ m/s the WT power generation increases with the underlying wind speed in region 2 ($R_{2}=[w_{\text{in}},w_{\text{Pmax}})$), until reaching its rated power generation of $P^{\text{W}}_{\max}=4200$ kW. In region 3 ($R_{3}=[w_{\text{Pmax}},w_{\text{out}})$) the power output remains constant on the rated maximal output. In order to prevent damage on the WT for high speeds, the generation will turn off at a cut-out speed $w_{\text{out}}=22.5$ m/s, summarized in region 4 ($R_{4}=[w_{\text{out}},\infty$). Note that the power curve takes constant values except for region 2, where the behavior has to be approximated by a function $P^{\text{W}}_{R_{2}}(w)$. Moreover, we choose a polynomial of degree 6 WANG2019109422 , to fit the underlying data points of the power curve in region 2, see Figure 11, i.e.,
$P^{\text{W}}_{R_{2}}(w)=a_{6}w^{6}+a_{5}w^{5}+a_{3}w^{4}+a_{3}w^{3}+a_{2}w^{2}+a_{1}w+a_{0},$ (158)
with coefficients $a_{0},a_{1},a_{2},a_{3},a_{4},a_{5},a_{6}\in\mathbb{R}$, which can be found in Table 3. We are aware that the behavior in region 2 can be mimicked by a cubic polynomial WANG2019109422 . However, our motivation for a polynomial of degree 6 is to ensure a smooth transition from the power output in region 2 to the rated power in region 3. To obtain a function that is a representation of the power curve, the following form is used